### v1

# Low complexity binary words avoiding $(5/2)^{+}$-powers

Abstract.
Rote words are infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$. Shallit and Shur, as well as Ollinger and Shallit, showed that there are Rote words that avoid $(5/2)^{+}$-powers and that this is best possible. In this note we give a structure theorem for the Rote words that avoid $(5/2)^{+}$-powers, confirming a conjecture of Ollinger and Shallit.
The work of James Currie is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number DDG-2024-00005].
The work of Narad Rampersad is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2019-04111].
1. Introduction
Two central concepts in combinatorics on words are power avoidance and factor complexity. Recently, Shallit and Shur [9] initiated the systematic investigation of the interplay between these two concepts. They examined two dual problems: 1) Given a particular power avoidance constraint, determine the range of possible factor complexities among all infinite words avoiding the specified power; and, 2) Given a class of words with specified factor complexities, determine the powers that are avoided by at least one word in this class.
A well-known classical result provides a solution for the latter problem for the class of Sturmian words, i.e., the class of infinite words that contain $n+1$ factors of length $n$ for every $n\geq 1$: the Fibonacci word avoids $(5+\sqrt{5})/2$-powers, and this is best possible among all Sturmian words [1]. The Sturmian words are the aperiodic infinite words with the least possible factor complexity function; Shallit and Shur [9] and Ollinger and Shallit [8] studied another class of infinite words with low complexity, namely, the Rote words, which are the infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$ [7]. Each paper gives an example of an infinite Rote word that avoids $(5/2)^{+}$-powers and shows that this is best possible among all Rote words. Ollinger and Shallit end their paper by observing that the Rote words that avoid $(5/2)^{+}$-powers appear to have a certain rigid structure reminiscent of the famous structure theorem of Restivo and Salemi [6] for the class of overlap-free words. In this note we show that this observation is indeed true and we obtain the precise structure theorem.
2. Preliminaries
For a positive integer $n$, let $\Sigma_{n}=\{0,1,\ldots,n-1\}$ and let $\Sigma_{n}^{*}$ denote the set of all finite words over $\Sigma_{n}$. By a binary word we mean a (finite or infinite) word over $\Sigma_{2}$. Let $w$ be a word and write $w=xyz$. Then the word $x$ is a prefix of $w$, the word $y$ is a factor of $w$, and the word $z$ is a suffix of $w$. A map $f:\Sigma_{m}^{*}\to\Sigma_{n}^{*}$ is a morphism if $f(xy)=f(x)f(y)$ for all $x$ and $y$.
Let $w$ have length $\ell$ and smallest period $p$. The exponent of $w$ is the quantity $k=\ell/p$ and $w$ is called a $k$-power. A $k^{+}$-power is a word with exponent $>k$. A $2$-power is a square and a $2^{+}$-power is an overlap. A word $w$ avoids $k$-powers (resp. $k^{+}$-powers) if none of its factors are $k^{\prime}$-powers for any $k^{\prime}>=k$ (resp. $k^{\prime}>k$); we also say that $w$ is $k$-power-free (resp. $k^{+}$-power-free).
Let ${\bf x}$ be an infinite word. The factor complexity of ${\bf x}$ is the function of $n$ that associates each length $n$ with the number of factors of ${\bf x}$ of length $n$. A Sturmian word is any infinite word with factor complexity $n+1$; a Rote word is any infinite word with factor complexity $2n$.
Definition 1.
Let $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ be the morphism given by
$g(0)$ $=011$
$g(1)$ $=0$
$g(2)$ $=01.$
Let $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be the morphism given by
$f(0)$ $=0121$
$f(1)$ $=021$
$f(2)$ $=01.$
Recently the morphisms $f$ and $g$ have proved useful in several constructions [3, 5, 8]. Dvořáková et al. showed that $g(f^{\omega}(0))$ avoids $(5/2)^{+}$-powers [5], while Ollinger and Shallit showed that the factor complexity of this word is $2n$ (i.e., that it is a Rote word) [8]. The latter authors conjectured that there is a structure theorem involving $f$ and $g$ for the class of $(5/2)^{+}$-power-free Rote words.
The prototypical structure theorem of this type was obtained by Restivo and Salemi [6] for the class of binary overlap-free words. In this case, the structure is specified using the Thue–Morse morphism
$\mu(0)$ $=01$
$\mu(1)$ $=10.$
Restivo and Salemi showed the following:
Theorem 1.
Let ${\boldsymbol{w}}\in\{0,1\}^{\omega}$ be overlap-free. Then a final segment of ${\boldsymbol{w}}$ has the form $\mu({\boldsymbol{u}})$ where ${\boldsymbol{u}}$ is overlap-free.
Karhumäki and Shallit [4] later showed that the same structure theorem holds for the class of binary $(7/3)^{+}$-power-free words. In this note we establish a similar structure theorem for the class of $(5/2)^{+}$-power-free Rote words in terms of the morphisms $f$ and $g$ given above. The reader may also compare the present structure theorem with the main result of [2], which establishes a structure theorem for the class of infinite $15/4$-power-free binary rich words (rich means that every factor of length $n$ contains $n$ distinct non-empty palindromes), and which also involves a sub-family of Rote words, namely the complementary symmetric Rote words.
3. Obtaining the structure theorem
Lemma 1.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Then both of the words $0110$ and $1001$ are factors of ${\boldsymbol{w}}$.
Proof.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $0110$ has length 14. Thus $0110$ (and symmetrically, $1001$) is a factor of ${\boldsymbol{w}}$. ∎
Lemma 2.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. At least 3 of the words in
$C=\{0010,0100,1011,1101\}$
are factors of ${\boldsymbol{w}}$.
Proof.
Six backtrack searches (one for each pair) show that the longest $\frac{5}{2}^{+}$ power free binary word omitting a pair of these factors has length 44. ∎
Lemma 3.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Both of $0011$ and $1100$ are factors of ${\boldsymbol{w}}$.
Proof.
Consider the set of seven binary words
$A=\{0010,0100,0101,1010,1011,1101,1100\}.$
For each word $a\in A$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing neither of $a$ and $0011$ as a factor has length no more than 31.
It follows that if $0011$ is not a factor of ${\boldsymbol{w}}$, then ${\boldsymbol{w}}$ contains the seven words of $A$ as length 4 factors. By Lemma 1, it also contains $0110$ and $1001$ as factors. However, now ${\boldsymbol{w}}$ contains 9 factors of length 4, contradicting the fact that its factor complexity is at most $2n$.
We conclude that $0011$ (and symmetrically, $1100$) is a factor of ${\boldsymbol{w}}$. ∎
Lemma 4.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. At least one of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$.
Proof.
Consider the set $D$ containing 17 binary words of length 9 given by
$D=\{00100110,01001100,10011001,00110010,01100100,11001001,$
$10010011,00110011,01100110,11001101,10011011,00110110,$
$101100 , 1$ 10110010 , 10110011 , 11001100 } .
For each word $d\in D$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing none of $d$, $0101$, and $1010$ as a factor has length no more than 88. It follows that if neither of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$, then every word of $D$ is a factor. However, this would imply that ${\boldsymbol{w}}$ contained 17 factors of length 8, contradicting the fact that its factor complexity is at most $2n$. We conclude that at least one of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$. ∎
Theorem 2.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Up to binary complement and/or reversal, the set of length 4 factors of ${\boldsymbol{w}}$ is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Proof.
By Lemma 1, the set of length 4 factors includes $0110$ and $1001$. By Lemma 3, the set of length 4 factors includes $0011$ and $1100$. Combining Lemmas 2 and 4 with the fact that ${\boldsymbol{w}}$ has at most 8 length 4 factors, the set of length 4 factors contains exactly 3 words from $C=\{0100,0010,1011,1101\}$ and exactly one word from $\{0101,1010\}$. Since each word of $C$ maps to each of the others under complement and/or reversal, assume without loss of generality that the 3 words from $C$ are $0010$, $0100$, and $1101$. Thus ${\boldsymbol{w}}$ does not contain the factor $1011$.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $1011$ or $1010$ as a factor has length 20. We conclude that ${\boldsymbol{w}}$ contains the factor $1010$, so that set of length 4 factors of ${\boldsymbol{w}}$ is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
For $u\in\Sigma_{3}^{*}$, denote the Parikh vector of $u$ by $\pi(u)$. Thus,
$\pi(u)=[|u|_{0},|u|_{1},|u|_{2}].$
For $x,y\in\Sigma_{3}^{*}$, we say that $\pi(x)>\pi(y)$ if:
  1. (1)
    We have $|x|_{i}\geq|y|_{i}$ for all $i\in\{0,1,2\}$.
  2. (2)
    For at least one $i\in\{0,1,2\}$ we have $|x|_{i}>|y|_{i}.$
Lemma 5.
Suppose that $u\in\Sigma_{3}^{*}$, $w\in\Sigma_{2}^{*}$, and $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is a non-erasing morphism. If $g(u)=w$, and $w$ avoids $\frac{5}{2}^{+}$ powers, then $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
Proof.
Suppose $u$ has a factor $xyxyx$ where $\pi(x)>\pi(y)$. Then $|g(x)|>|g(y)|$, and $w$ contains the $\frac{5}{2}^{+}$ power $g(x)g(y)g(x)g(y)g(x)$. ∎
Lemma 6.
Suppose that $u,v\in\Sigma_{3}^{*}$, and $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is a non-erasing morphism. If $f(v)=u$, and $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$, then $v$ has no factor $XYXYX$ where $\pi(X)>\pi(Y)$.
Proof.
Suppose $v$ has a factor $XYXYX$ where $\pi(X)>\pi(Y)$. Let $x=f(X)$ and $y=f(Y)$. Then $\pi(x)>\pi(y)$, and $u$ contains the factor $xyxyx$. ∎
Definition 2.
Call a word $u\in\Sigma_{3}^{*}$ proper if:
  1. (1)
    Word $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
  2. (2)
    None of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of $u$.
Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ proper if all of its finite factors are proper.
Lemma 7.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of ${\boldsymbol{w}}$ is
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Then a final segment of ${\boldsymbol{w}}$ has the form $g({\boldsymbol{u}})$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
Proof.
Since $111$ is not a factor of ${\boldsymbol{w}}$, any final segment of ${\boldsymbol{w}}$ beginning with 00 has the form $g({\boldsymbol{u}})$ for some ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$. We will show, replacing ${\boldsymbol{u}}$ by one of its final segments if necessary, that ${\boldsymbol{u}}$ is proper.
The fact that ${\boldsymbol{u}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 5.
We conclude by showing that none of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of ${\boldsymbol{u}}$:
Word $00$: If $00$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(00)=011011$. However, then ${\boldsymbol{w}}$ contains $1011$, which is not in $F$.
Word $11$: If $11$ is a factor of ${\boldsymbol{u}}$, then $11a$ is a factor of ${\boldsymbol{u}}$ for some $a\in\Sigma_{3}$. Then ${\boldsymbol{w}}$ contains the factor $g(11a)$, which starts with $g(1)g(1)0=000$, a $\frac{5}{2}^{+}$ power.
Word $22$: If $22$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(22)=0101\notin F$.
Word $20$: If $20$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(20)=01011$, which starts with $0101\notin F$.
Word $10101$: If $10101$ is not a factor of ${\boldsymbol{u}}$ more than once, replace ${\boldsymbol{u}}$ by one of its final segments not containing $10101$.
On the other hand, if $10101$ is a factor of ${\boldsymbol{u}}$ more than once, then $a10101b$ is a factor of ${\boldsymbol{u}}$ for some $a,b\in\Sigma_{3}$. Since $11$ is not a factor of ${\boldsymbol{u}}$, we can in fact specify that $a\in\Sigma_{3}-\{1\}.$ This implies that $1$ is the last letter of $g(a)$. Also, 00 is the first letter of $g(b)$. Then ${\boldsymbol{w}}$ contains the factor $g(a10101b)$, which contains $100110011001g(10101)0=$ 10011001100, a $\frac{5}{2}^{+}$ power.
Word $2121$: If $2121$ is a factor of ${\boldsymbol{u}}$, then $2121a$ is a factor of ${\boldsymbol{u}}$ for some $a\in\Sigma_{3}-\{1\}$. Thus $01$ is a prefix of $g(a)$, and ${\boldsymbol{w}}$ contains the factor $g(2121)01=01001001$, a $\frac{5}{2}^{+}$ power.
[…]
Lemma 8.
Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be proper. Then a final segment of ${\boldsymbol{u}}$ has the form $f({\boldsymbol{v}})$ for some proper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
Proof.
Assume without loss of generality, replacing ${\boldsymbol{u}}$ by a final segment if necessary, that ${\boldsymbol{u}}$ starts with the letter 00. Write ${\boldsymbol{u}}$ as a concatenation of 00-blocks, i.e., words which start with 00, and contain the letter 00 exactly once. Since ${\boldsymbol{u}}$ is proper, it does not contain a factor $00$ or $20$. It follows that its 00-blocks end with $1$. Since every occurrence of $2$ in a proper word can only be followed by a $1$, ${\boldsymbol{u}}$ cannot contain the factor $212$; otherwise it would contain the forbidden factor $2121$. We conclude that the 00-blocks of ${\boldsymbol{u}}$ are among $01$, $021$ and $0121$. Notice that the last two of these words are always followed by a 00 in ${\boldsymbol{u}}$. We thus conclude that a final segment of ${\boldsymbol{u}}$ has the form $f({\boldsymbol{v}})$ for some ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$. We will show, replacing ${\boldsymbol{v}}$ by one of its final segments if necessary, that ${\boldsymbol{v}}$ is proper.
The fact that ${\boldsymbol{v}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 6.
We conclude by showing that a final segment of ${\boldsymbol{v}}$ contains none of the words $22$, $20$, $00$, $11$, $10101$, $2121$, or $10210210$ as a factor:
Word $22$: If $22$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $22$. Otherwise, $22$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $101011f(22)=$ 10101 . This is impossible, since ${\boldsymbol{w}}$ is proper.
Word $20$: If $20$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $20$. Otherwise, $20$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $10101211f(20)=$ 1010121, which starts with $10101$. This is impossible, since ${\boldsymbol{w}}$ is proper.
Word $00$: Suppose that $00$ is a factor of ${\boldsymbol{v}}$. If $000$ is a factor of ${\boldsymbol{v}}$, then $f(000)=012101210121$ is a factor of ${\boldsymbol{u}}$. However $012101210121=xyxyx$ where $x=0121$, $y=\epsilon$, and cannot be a factor of ${\boldsymbol{u}}$. It follows that $000$ is not a factor of ${\boldsymbol{v}}$.
If $00$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $00$. Otherwise, $00$ is a factor of ${\boldsymbol{v}}$ more than once, so that $100a$ is a factor of ${\boldsymbol{v}}$ for some $a\in\Sigma_{3}$. Since $f(a)$ starts with 00, this implies that ${\boldsymbol{u}}$ has a factor $f(100a)$, starting with $021012101210$. This contains the word $xyxyx$ where $x=210$, $y=1$, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $11$: If $11$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $11$. Otherwise, $11$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $102102101f(11)0=$ 10210210, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $10101$: If $10101$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $10101$. Otherwise, $10101$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $1f(10101)0=1021012102101210210=xyxyx$, where $x=10210$ and $y=12$. This is impossible, since ${\boldsymbol{u}}$ is proper.
Word $2121$: If $2121$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $2121$. Otherwise, $2121$ is a factor of ${\boldsymbol{v}}$ more than once. Since $22$ is not a factor of $v$, $a2121b$ is a factor of ${\boldsymbol{v}}$ for some $a,b\in\Sigma_{3}$ where $a\neq 2$. Then $f(a)$ ends in $21$ and $f(b)$ begins with 00. Thus ${\boldsymbol{u}}$ contains the factor $21f(2121)0=2101021010210=xyxyx$ where $x=210$, and $y=10$, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $10210210$: If $10210210$ is a factor of ${\boldsymbol{v}}$, then ${\boldsymbol{u}}$ contains the factor $f(10210210)=0210121010210121010210121=xyxyx$, where $x=0210121$, and $y=01$, which is impossible, since ${\boldsymbol{u}}$ is proper. ∎
Theorem 3.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of ${\boldsymbol{w}}$ is
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Then for every positive integer $n$, a final segment of ${\boldsymbol{w}}$ has the form $g(f^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
Proof.
This follows from Lemmas 8 and 7 by induction. ∎
Many open problems remain concerning the relationship between low factor complexity and avoidable powers. As an example, moving to a ternary alphabet, Shallit and Shur [9] showed that the word
[…]
generated by the morphism $0\to 01,1\to 2,2\to 02$, has critical exponent $2.4808627\cdots$ and factor complexity $12n+$ 1 for all $n\geq 1$. They conjecture that this exponent is minimal among all infinite ternary words with complexity $12n+$ 1. This conjecture remains open. If confirmed, it would be natural to explore the possibility of a structure theorem for this class of words.