### v3

# Low complexity binary words avoiding $(5/2)^{+}$-powers

Abstract
Rote words are infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$. Shallit and Shur, as well as Ollinger and Shallit, showed that there are Rote words that avoid $(5/2)^{+}$-powers and that this is best possible. In this note we give a structure theorem for the Rote words that avoid $(5/2)^{+}$-powers, confirming a conjecture of Ollinger and Shallit.
keywords:
Rote word, factor complexity, $(5/2)^{+}$-power, structure theorem
1 Introduction
Two central concepts in combinatorics on words are power avoidance and factor complexity. Recently, Shallit and Shur (2019) initiated the systematic investigation of the interplay between these two concepts. They examined two dual problems: 1) Given a particular power avoidance constraint, determine the range of possible factor complexities among all infinite words avoiding the specified power; and, 2) Given a class of words with specified factor complexities, determine the powers that are avoided by at least one word in this class.
A well-known classical result provides a solution for the latter problem for the class of Sturmian words, i.e., the class of infinite words that contain $n+1$ factors of length $n$ for every $n\geq 1$: the Fibonacci word avoids $(5+\sqrt{5})/2$-powers, and this is best possible among all Sturmian words (Carpi and de Luca (2000)). The Sturmian words are the aperiodic infinite words with the least possible factor complexity function; Shallit and Shur (2019) and Ollinger and Shallit (2024) studied another class of infinite words with low complexity, namely, the Rote words, which are the infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$ (Rote (1994)). Each paper gives an example of an infinite Rote word that avoids $(5/2)^{+}$-powers and shows that this is best possible among all Rote words. Ollinger and Shallit end their paper by observing that the Rote words that avoid $(5/2)^{+}$-powers appear to have a certain rigid structure reminiscent of the famous structure theorem of Restivo and Salemi (1985) for the class of overlap-free words. In this note we obtain the precise structure theorem.
To state our first structure theorem, we introduce proper words and antiproper words, which are ternary words.
Definition 1.
For $u\in\Sigma_{3}^{*}$, denote the Parikh vector of $u$ by $\pi(u)$so that $\pi(u)=[|u|_{0},|u|_{1},|u|_{2}].$ For $x,y\in\Sigma_{3}^{*}$, we say that $\pi(x)>\pi(y)$ if:
  1. 1.
    We have $|x|_{i}\geq|y|_{i}$ for all $i\in\Sigma_{3}$.
  2. 2.
    For at least one $i\in\Sigma_{3}$ we have $|x|_{i}>|y|_{i}.$
Call a word $u\in\Sigma_{3}^{*}$ proper if:
  1. 1.
    Word $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$.
  2. 2.
    None of the words $00$, $11$, $22$, 2020, $10101$, $2121$, or $10210210$ is a factor of $u$.
Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ proper if all of its finite factors are proper.
Call a word $u\in\Sigma_{3}^{*}$ antiproper if its reverse $u^{R}$ is proper. Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ antiproper if all of its finite factors are antiproper.
Proper words obey a structure theorem similar to that of Restivo and Salemi, as do antiproper words. We introduce a morphism $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ and its reverse $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$, given by:
$f(0)$ $=0121$ $h(0)$ $=1210$
$f(1)$ $=021$ $h(1)$ $=120$
$f(2)$ $=01$ $h(2)$ $=10.$
Theorem 1.
(First Structure Theorem)
  1. 1.
    Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be proper. Then a final segment of $\boldsymbol{u}$} has the form $f({\boldsymbol{v}})$ for some proper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
  2. 2.
    Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be antiproper. Then a final segment of $\boldsymbol{u}$} has the form $h({\boldsymbol{v}})$ for some antiproper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
For our second structure theorem we consider the length-4 factors of a Rote word.
Definition 2.
Let $F$ be the set
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Let $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ be the morphism given by
$g(0)$ $=011$
$g(1)$ $=0$
$g(2)$ $=01.$
We denote the complement of a binary word $w$ by $\overline{w}$; thus $\overline{1101}=0010$. We extend this notation to binary languages in the usual way. We denote the reversal of a word $w$ by $w^{R}$; thus $1101^{R}=1011$. We extend this notation to languages in the usual way.
We can now characterize the structure of Rote words that avoid $(5/2)^{+}$-powers.
Theorem 2.
(Second Structure Theorem) Let $\boldsymbol{w}$} be a Rote word that avoids $(5/2)^{+}$-powers. There are four cases:
  1. 1.
    The set of length-4 factors of $\boldsymbol{w}$} is $F$. For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $g(f^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
  2. 2.
    The set of length-4 factors of $\boldsymbol{w}$} is $\bar{F}$. For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $\overline{g(f^{n}({\boldsymbol{u}}))}$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
  3. 3.
    The set of length-4 factors of $\boldsymbol{w}$} is $F^{R}$. For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $g(h^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
  4. 4.
    The set of length-4 factors of $\boldsymbol{w}$} is $\overline{F^{R}}$. For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $\overline{g(h^{n}({\boldsymbol{u}}))}$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
2 Preliminaries
For a positive integer $n$, let $\Sigma_{n}=\{0,1,\ldots,n-1\}$ and let $\Sigma_{n}^{*}$ denote the set of all finite words over $\Sigma_{n}$. By a binary word we mean a (finite or infinite) word over $\Sigma_{2}$. Let $w$ be a word and write $w=xyz$. Then the word $x$ is a prefix of $w$, the word $y$ is a factor of $w$, and the word $z$ is a suffix of $w$. A map $f:\Sigma_{m}^{*}\to\Sigma_{n}^{*}$ is a morphism if $f(xy)=f(x)f(y)$ for all $x$ and $y$.
Let $w$ have length $\ell$ and smallest period $p$. The exponent of $w$ is the quantity $k=\ell/p$ and $w$ is called a $k$-power. A $k^{+}$-power is a word with exponent >k>k. A $2$-power is a square and a $2^{+}$-power is an overlap. A word $w$ avoids $k$-powers (resp. $k^{+}$-powers) if none of its factors are $k^{\prime}$-powers for any $k^{\prime}>=k$ (resp. $k^{\prime}>k$); we also say that $w$ is $k$-power-free (resp. $k^{+}$-power-free).
Let $\bf x$} be an infinite word. The factor complexity of $\bf x$} is the function of $n$ that associates each length $n$ with the number of factors of $\bf x$} of length $n$. A Sturmian word is any infinite word with factor complexity $n+1$; a Rote word is any infinite word with factor complexity $2n$.
Recently the morphisms $f$ and $g$ have proved useful in several constructions (Currie et al. (2023); Dvořáková et al. (2024); Ollinger and Shallit (2024)). Dvořáková et al. (2024) showed that $g(f^{\omega}(0))$ avoids $(5/2)^{+}$-powers, while Ollinger and Shallit (2024) showed that the factor complexity of this word is $2n$ (i.e., that it is a Rote word). The latter authors conjectured that there is a structure theorem involving $f$ and $g$ for the class of $(5/2)^{+}$-power-free Rote words.
A prototypical structure theorem of this type was obtained by Restivo and Salemi (1985) for the class of binary overlap-free words. (See also Fife (1980); Shur (1996); Shallit (2011).) In this case, the structure is specified using the Thue–Morse morphism
$\mu(0)$ $=01$
$\mu(1)$ $=10.$
Restivo and Salemi showed the following:
Theorem 3.
Let ${\boldsymbol{w}}\in\Sigma_{2}^{\omega}$ be overlap-free. Then a final segment of $\boldsymbol{w}$} has the form $\mu({\boldsymbol{u}})$ where $\boldsymbol{u}$} is overlap-free.
Karhumäki and Shallit (2004) later showed that the same structure theorem holds for the class of binary $(7/3)^{+}$-power-free words. In this note we establish a similar structure theorem for the class of $(5/2)^{+}$-power-free Rote words in terms of the morphisms $f$ and $g$ given above. The reader may also compare the present structure theorem with the main result of Currie et al. (2020–2021), which establishes a structure theorem for the class of infinite $14/5$-power-free binary rich words (rich means that every factor of length $n$ contains $n$ distinct non-empty palindromes), and which also involves a sub-family of Rote words, namely the complementary symmetric Rote words.
3 Obtaining the structure theorem
Lemma 1.
Let $\boldsymbol{w}$} be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Then both of the words $0110$ and $1001$ are factors of $\boldsymbol{w}$}.
Proof.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $0110$ has length 14. Thus $0110$ (and symmetrically, $1001$) is a factor of $\boldsymbol{w}$}. ∎
Lemma 2.
Let $\boldsymbol{w}$} be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. At least 3 of the words in
$C=\{0010,0100,1011,1101\}$
are factors of $\boldsymbol{w}$}.
Proof.
Six backtrack searches (one for each pair) show that the longest $\frac{5}{2}^{+}$ power free binary word omitting a pair of these factors has length 44. ∎
Lemma 3.
Let $\boldsymbol{w}$} be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Both of $0011$ and $1100$ are factors of $\boldsymbol{w}$}.
Proof.
Consider the set of seven binary words
$A=\{0010,0100,0101,1010,1011,1101,1100\}.$
For each word $a\in A$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing neither of $a$ and $0011$ as a factor has length no more than 31.
It follows that if $0011$ is not a factor of $\boldsymbol{w}$}, then $\boldsymbol{w}$} contains the seven words of $A$ as length 4 factors. By Lemma 1, it also contains $0110$ and $1001$ as factors. However, now $\boldsymbol{w}$} contains 9 factors of length 4, contradicting the fact that its factor complexity is at most $2n$.
We conclude that $0011$ (and symmetrically, $1100$) is a factor of $\boldsymbol{w}$}. ∎
Lemma 4.
Let $\boldsymbol{w}$} be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. At least one of $0101$ and $1010$ is a factor of $\boldsymbol{w}$}.
Proof.
Consider the set $D$ containing 17 binary words of length 9 given by
$D=\{00100110,01001100,10011001,00110010,01100100,11001001,$
10010011,00110011,01100110,11001101,10011011,00110110,10010011,00110011,01100110,11001101,10011011,00110110,
01101100,11011001,10110010,10110011,11001100}.01101100,11011001,10110010,10110011,11001100\}.
For each word $d\in D$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing none of $d$, $0101$, and $1010$ as a factor has length no more than 88. It follows that if neither of $0101$ and $1010$ is a factor of $\boldsymbol{w}$}, then every word of $D$ is a factor. However, this would imply that $\boldsymbol{w}$} contained 17 factors of length 8, contradicting the fact that its factor complexity is at most $2n$. We conclude that at least one of $0101$ and $1010$ is a factor of $\boldsymbol{w}$}. ∎
Theorem 4.
Let $\boldsymbol{w}$} be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Up to binary complement and/or reversal, the set of length 4 factors of $\boldsymbol{w}$} is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Proof.
By Lemma 1, the set of length 4 factors includes $0110$ and $1001$. By Lemma 3, the set of length 4 factors includes $0011$ and $1100$. Combining Lemmas 2 and 4 with the fact that $\boldsymbol{w}$} has at most 8 length 4 factors, the set of length 4 factors contains exactly 3 words from $C=\{0100,0010,1011,1101\}$ and exactly one word from $\{0101,1010\}$. Since each word of $C$ maps to each of the others under complement and/or reversal, assume without loss of generality that the 3 words from $C$ are $0010$, $0100$, and $1101$. Thus $\boldsymbol{w}$} does not contain the factor $1011$.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $1011$ or $1010$ as a factor has length 20. We conclude that $\boldsymbol{w}$} contains the factor $1010$, so that set of length 4 factors of $\boldsymbol{w}$} is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Lemma 5.
Suppose that $u\in\Sigma_{3}^{*}$, $w\in\Sigma_{2}^{*}$, and $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is a non-erasing morphism. If $g(u)=w$, and $w$ avoids $\frac{5}{2}^{+}$ powers, then $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$.
Proof.
Suppose $u$ has a factor xyxyxxyxyx where $\pi(x)>\pi(y)$. Then $|g(x)|>|g(y)|$, and $w$ contains the $\frac{5}{2}^{+}$ power $g(x)g(y)g(x)g(y)g(x)$. ∎
Lemma 6.
Suppose that $u,v\in\Sigma_{3}^{*}$, and $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is a non-erasing morphism. If $f(v)=u$, and $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$, then $v$ has no factor $XYXYX$ where $\pi(X)>\pi(Y)$.
Proof.
Suppose $v$ has a factor $XYXYX$ where $\pi(X)>\pi(Y)$. Let $x=f(X)$ and $y=f(Y)$. Then $\pi(x)>\pi(y)$, and $u$ contains the factor xyxyxxyxyx. ∎
Lemma 7.
Let $\boldsymbol{w}$} be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of $\boldsymbol{w}$} is
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Then a final segment of $\boldsymbol{w}$} has the form $g({\boldsymbol{u}})$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
Proof.
Since $111$ is not a factor of $\boldsymbol{w}$}, any final segment of $\boldsymbol{w}$} beginning with 0 has the form $g({\boldsymbol{u}})$ for some ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$. We will show, replacing $\boldsymbol{u}$} by one of its final segments if necessary, that $\boldsymbol{u}$} is proper.
The fact that $\boldsymbol{u}$} has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$ follows from Lemma 5.
We conclude by showing that none of the words $00$, $11$, $22$, 2020, $10101$, $2121$, or $10210210$ is a factor of $\boldsymbol{u}$}:
Word $00$: If $00$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor $g(00)=011011$. However, then $\boldsymbol{w}$} contains $1011$, which is not in $F$.
Word $11$: If $11$ is a factor of $\boldsymbol{u}$}, then $11a$ is a factor of $\boldsymbol{u}$} for some $a\in\Sigma_{3}$. Then $\boldsymbol{w}$} contains the factor $g(11a)$, which starts with $g(1)g(1)0=000$, a $\frac{5}{2}^{+}$ power.
Word $22$: If $22$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor $g(22)=0101\notin F$.
Word $20$: If $20$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor $g(20)=01011$, which starts with $0101\notin F$.
Word $10101$: If $10101$ is not a factor of $\boldsymbol{u}$} more than once, replace $\boldsymbol{u}$} by one of its final segments not containing $10101$.
On the other hand, if $10101$ is a factor of $\boldsymbol{u}$} more than once, then $a10101b$ is a factor of $\boldsymbol{u}$} for some $a,b\in\Sigma_{3}$. Since $11$ is not a factor of $\boldsymbol{u}$}, we can in fact specify that $a\in\Sigma_{3}-\{1\}.$ This implies that $1$ is the last letter of $g(a)$. Also, 0 is the first letter of $g(b)$. Then $\boldsymbol{w}$} contains the factor $g(a10101b)$, which contains $1g(10101)0=10011001100$, a $\frac{5}{2}^{+}$ power.
Word $2121$: If $2121$ is a factor of $\boldsymbol{u}$}, then $2121a$ is a factor of $\boldsymbol{u}$} for some $a\in\Sigma_{3}-\{1\}$. Thus $01$ is a prefix of $g(a)$, and $\boldsymbol{w}$} contains the factor $g(2121)01=01001001$, a $\frac{5}{2}^{+}$ power.
Word $10210210$: If $10210210$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor
$g(10210210)=0011010011010011,$
a $\frac{5}{2}^{+}$ power. ∎
We can now prove Theorem 1:
of Theorem 1.
Assume without loss of generality, replacing $\boldsymbol{u}$} by a final segment if necessary, that $\boldsymbol{u}$} starts with the letter 0. Write $\boldsymbol{u}$} as a concatenation of 0-blocks, i.e., words which start with 0, and contain the letter 0 exactly once. Since $\boldsymbol{u}$} is proper, it does not contain a factor $00$ or $20$. It follows that its 0-blocks end with $1$. Since every occurrence of $2$ in a proper word can only be followed by a $1$, $\boldsymbol{u}$} cannot contain the factor $212$; otherwise it would contain the forbidden factor $2121$. We conclude that the 0-blocks of $\boldsymbol{u}$} are among $01$, $021$ and $0121$. Notice that the last two of these words are always followed by a 0 in $\boldsymbol{u}$}. We thus conclude that a final segment of $\boldsymbol{u}$} has the form $f({\boldsymbol{v}})$ for some ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$. We will show, replacing $\boldsymbol{v}$} by one of its final segments if necessary, that $\boldsymbol{v}$} is proper.
The fact that $\boldsymbol{v}$} has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$ follows from Lemma 6.
We conclude by showing that a final segment of $\boldsymbol{v}$} contains none of the words $22$, 2020, $00$, $11$, $10101$, $2121$, or $10210210$ as a factor:
Word $22$: If $22$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $22$. Otherwise, $22$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(22)=10101.$ This is impossible, since $\boldsymbol{u}$} is proper.
Word $20$: If $20$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $20$. Otherwise, $20$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(20)=1010121$, which starts with $10101$. This is impossible, since $\boldsymbol{u}$} is proper.
Word $00$: Suppose that $00$ is a factor of $\boldsymbol{v}$}. If $000$ is a factor of $\boldsymbol{v}$}, then $f(000)=012101210121$ is a factor of $\boldsymbol{u}$}. However $012101210121=xyxyx$ where $x=0121$, $y=\epsilon$, and cannot be a factor of $\boldsymbol{u}$}. It follows that $000$ is not a factor of $\boldsymbol{v}$}.
If $00$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $00$. Otherwise, $00$ is a factor of $\boldsymbol{v}$} more than once, so that $100a$ is a factor of $\boldsymbol{v}$} for some $a\in\Sigma_{3}$. Since $f(a)$ starts with 0, this implies that $\boldsymbol{u}$} has a factor $f(100a)$, starting with $021012101210$. This contains the word xyxyxxyxyx where $x=210$, $y=1$, which is impossible, since $\boldsymbol{u}$} is proper.
Word $11$: If $11$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $11$. Otherwise, $11$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(11)0=10210210$, which is impossible, since $\boldsymbol{u}$} is proper.
Word $10101$: If $10101$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $10101$. Otherwise, $10101$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(10101)0=1021012102101210210=xyxyx$, where $x=10210$ and $y=12$. This is impossible, since $\boldsymbol{u}$} is proper.
Word $2121$: If $2121$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $2121$. Otherwise, $2121$ is a factor of $\boldsymbol{v}$} more than once. Since $22$ is not a factor of $v$, $a2121b$ is a factor of $\boldsymbol{v}$} for some $a,b\in\Sigma_{3}$ where $a\neq 2$. Then $f(a)$ ends in $21$ and $f(b)$ begins with 0. Thus $\boldsymbol{u}$} contains the factor $21f(2121)0=2101021010210=xyxyx$ where $x=210$, and $y=10$, which is impossible, since $\boldsymbol{u}$} is proper.
Word $10210210$: If $10210210$ is a factor of $\boldsymbol{v}$}, then $\boldsymbol{u}$} contains the factor
f(10210210)=0210121010210121010210121=xyxyx,f(10210210)=0210121010210121010210121=xyxyx,
where $x=0210121$, and $y=01$, which is impossible, since $\boldsymbol{u}$} is proper.
The proof for antiproper words is the same, mutatis mutandi. ∎
We can now prove Theorem 2:
of Theorem 2.
The first case follows from Theorem 1 and Lemma 7 by induction. The other cases follow, mutatis mutandi. ∎
As an example of the third case of the theorem, Shallit and Shur (2019) consider a word $\tau({\bf G})$, where $\tau:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is the morphism given by
$\tau(0)$ $=0$
$\tau(1)$ $=01$
$\tau(2)$ $=011$
and $\bf G$} is the fixed point of $\theta$, where $\theta:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is the morphism given by
$\theta(0)$ $=01$
$\theta(1)$ $=2$
$\theta(2)$ $=02.$
Letting $\sigma$ be the permutation $0\to 1\to 2\to 0$, one checks that $\tau=g\sigma$ and $\theta^{2}=\sigma^{-1}h\sigma$. It follows that
$\tau({\bf G})$ $=\tau(\theta^{2n}(\theta^{\omega}(0)))$
$=g\sigma((\sigma^{-1}h\sigma)^{n-1}(\sigma^{-1}h\sigma(\theta^{\omega}(0)))$
$=gh^{n-1}({\boldsymbol{u}})),$
where ${\boldsymbol{u}}=h(\sigma(\theta^{\omega}(0)))$, which is antiproper.
Many open problems remain concerning the relationship between low factor complexity and avoidable powers. Moving to a ternary alphabet, Shallit and Shur (2019) showed that the word
${\bf G}=0120201020120102012\cdots$
has critical exponent $2.4808627\cdots$ and factor complexity $2n+1$ for all $n\geq 1$. They conjectured that this exponent is minimal among all infinite ternary words with complexity $2n+1$. This conjecture was recently confirmed by Currie (2025). It would thus be natural to explore the possibility of a structure theorem for this class of words.
4 Appendix: Python code and output
The backtrack searches mentioned in the paper run quickly in Python. Here is our code and its output:
def fhpf(w): #Word w is 5/2+ power suffix free
    p=1
    while (5*p<2*len(w)):
        if (w[(-(p+1)//2)-p:]==w[(-(p+1)//2)-2*p:-p]):
            return(False)
        p=p+1
    return(True)
def good(w): # Word w has no suffix which is a 5/2+ power, or is in the
             # set Factors.
    for f in Factors:
        k=len(f)
        if ((len(w)>= k)and(w[-k:]==f)):
            return(False)
    return(fhpf(w))
def search(target): # This returns the lexicographically least word not
                    # containing a 5/2+ power or a word in the set Factors
    w=’’
    Max=0
    while (len(w)<=target):
        if (good(w)):
            Max=max(Max,len(w))
            if (len(w)==target):
                return(w)
            w+=’0’
        else:
            while((len(w)>0) and (w[-1]==’1’)):
                w=w[:-1]
            if(len(w)==0):
                print(’Longest 5/2+-power-free word with no factor in ’,Factors,’
                    has length ’,Max)
                return()
            c=chr(ord(w[-1])+1)
            w=w[:-1]
            w+=c
    return()
# Lemma 1
print(’=============’)
print(’Computations for Lemma 1’)
print(’ ’)
Factors=[’0110’]
search(200)
# Lemma 2
print(’ ’)
print(’=============’)
print(’Computations for Lemma 2’)
print(’ ’)
C=[’0010’,’0100’,’1011’,’1101’]
for i in range(4):
    for j in range(i,4):
        Factors=[C[i]]
        Factors.append(C[j])
        search(200)
# Lemma 3
print(’ ’)
print(’=============’)
print(’Computations for Lemma 3’)
print(’ ’)
A=[’0010’,’0100’,’0101’,’1010’,’1011’,’1101’,’1100’]
for j in A:
    Factors=[’0011’]
    Factors.append(j)
    search(200)
# Lemma 4
print(’ ’)
print(’=============’)
print(’Computations for Lemma 4’)
print(’ ’)
D=[’00100110’,’01001100’,’10011001’,’00110010’,
 ’01100100’,’11001001’,’10010011’,’$0011$’,
 ’01100110’,’11001101’,’10011011’,’00110110’,
 ’01101100’,’11011001’,’10110010’,’10110011’,’$1100$’]
for j in D:
    Factors=[’0101’,’$10$’]
    Factors.append(j)
    search(200)
# Theorem 4
print(’ ’)
print(’=============’)
print(’Computations for Theorem 4’)
print(’ ’)
Factors=[’1011’,’$10$’]
search(200)
Computations for Lemma 1
Longest 5/2+-power-free word with no factor in  [’0110’]  has length  14
Computations for Lemma 2
Longest 5/2+-power-free word with no factor in  [’0010’, ’0100’]  has length  44
Longest 5/2+-power-free word with no factor in  [’0010’, ’1011’]  has length  28
Longest 5/2+-power-free word with no factor in  [’0010’, ’1101’]  has length  13
Longest 5/2+-power-free word with no factor in  [’0100’, ’1011’]  has length  13
Longest 5/2+-power-free word with no factor in  [’0100’, ’1101’]  has length  28
Longest 5/2+-power-free word with no factor in  [’1011’, ’1101’]  has length  44
Computations for Lemma 3
Longest 5/2+-power-free word with no factor in  [’0011’, ’0010’]  has length  15
Longest 5/2+-power-free word with no factor in  [’0011’, ’0100’]  has length  31
Longest 5/2+-power-free word with no factor in  [’0011’, ’0101’]  has length  12
Longest 5/2+-power-free word with no factor in  [’0011’, ’1010’]  has length  18
Longest 5/2+-power-free word with no factor in  [’0011’, ’1011’]  has length  15
Longest 5/2+-power-free word with no factor in  [’0011’, ’1101’]  has length  31
Longest 5/2+-power-free word with no factor in  [’0011’, ’1100’]  has length  30
Computations for Lemma 4
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00100110’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01001100’]  has length  50
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10011001’]  has length  33
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110010’]  has length  50
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01100100’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001001’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10010011’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110011’]  has length  52
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01100110’]  has length  33
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001101’]  has length  50
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10011011’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110110’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01101100’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11011001’]  has length  24
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10110010’]  has length  88
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10110011’]  has length  50
Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001100’]  has length  52
Computations for Theorem 4
Longest 5/2+-power-free word with no factor in  [’1011’, ’1010’]  has length  20